OpenStudy (anonymous):

An equation is shown below: 2x + y = 3 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@StudyGurl14 @myininaya

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (campbell_st):

part A ... graph the line... which shows all ordered pairs that form the line or rewrite the line in slope intercept form part B substitute 3 values of x to find 3 values of y... say, x = 0, x = 1, x = 2 hope that helps

OpenStudy (anonymous):

what do you mean rewrite the line in slope intercept form rewrite the equations?

OpenStudy (anonymous):

i need help on part c @campbell_st

OpenStudy (anonymous):

@AriPotta

OpenStudy (campbell_st):

I mean make y the subject of the equation slope intercept form is y = mx + b m = slope b = y-intercept

OpenStudy (anonymous):

okay but i did that so i need help on part 3 now

OpenStudy (anonymous):

Ari can you help

OpenStudy (aripotta):

make a line the is parallel to the given line

OpenStudy (anonymous):

y=2x-3

OpenStudy (campbell_st):

ok... so what is the slope of your line...?

OpenStudy (anonymous):

2

OpenStudy (aripotta):

but once you convert 2x + y = 3 into slope-intercept form, is y = -2x + 3 not y-= 2x - 3

OpenStudy (aripotta):

not y = 2x - 3*

OpenStudy (anonymous):

oh okay so then what

OpenStudy (aripotta):

so parallel lines have the same slope, but different y-intercepts. so just change - 3 to something else and put it in standard form

OpenStudy (anonymous):

wait what do you mean

OpenStudy (anonymous):

so it could be y=-2x+2?

OpenStudy (aripotta):

yea that could work. put it in standard form and maybe multiply the whole equation by like 3 or something to change it up a bit

OpenStudy (anonymous):

so now it would be 2x+y=3 and 2x+4=3?

OpenStudy (aripotta):

2x + y = 3 and 2x + y = 2

OpenStudy (anonymous):

okay thanks!

OpenStudy (aripotta):

no problemo :)